OpenStudy (ny,ny):

HELP is the ordered pair on the graph of the given equation? (-6,5); y=1/2x+2

OpenStudy (mertsj):

Did you plug it in to see if it makes the sentence true?

OpenStudy (ny,ny):

plug it in for what?

OpenStudy (mertsj):

Do you see the ordered pair (-6,5) in your question?

OpenStudy (ny,ny):

there's no graph.

OpenStudy (anonymous):

You have to determine if when x = -6 and y = 5, the equation y=1/2x+2 if true. To do this, plug in -6 instead of x and 5 instead of y in the equation.

OpenStudy (ny,ny):

Ohh yea. 5=1/2(-6)+2 5=3+2 5=5 right?

OpenStudy (anonymous):

That's exactly it !

OpenStudy (ny,ny):

yay thanks.

OpenStudy (ny,ny):

=)